# Efficiently Computing Equilibria in Budget-Aggregation Games

### v2

@@ 8 Weighted agents
In practice, one might want to assign different shares to the agents, e.g., when they represent groups of citizens of different size, or citizens in different tax brackets.777 The weighted setting was studied by Brandt et al. (2023, 2025) in the context of donor coordination. They proved that, with Leontief utilities, the unique equilibrium distribution is the unique distribution that maximizes the weighted product of agents’ utilities.
We define a weighted budget-aggregation game as a tuple $(N,A,(u_{i})_{i\in N},(w_{i})_{i\in N})$, where $N$, AA and $(u_{i})_{i\in N}$ are as in Definition 3.1, and $w_{i}$ is the weight of agent $i$, with $\sum_{i\in N}w_{i}=1$. The strategy set of each agent $i$ is given by
@@
If all weights are rational numbers, and the utility functions satisfy some concavity conditions, then the weighted setting can be reduced to the unweighted setting by agent cloning, resulting in a pseudo-polynomial time algorithm. We formalize the agent cloning process below.
@@
Given a weighted budget-aggregation game $(N,A,(u_{i})_{i\in N},(w_{i})_{i\in N})$ where all $w_{i}$ are rational numbers, let $D$ be their least common denominator, such that $w_{i}=q_{i}/D$ for all $i\in N$, and all $q_{i}$ are integers.
The corresponding cloned game is an (unweighted) budget-aggregation game in which the set of agents $N^{c}$ contains $q_{i}$ clones of each agent $i$ (with the same utility function $u_{i}$), so that $|N^{c}|=\sum_{i\in N}q_{i}=D$.
The concavity condition we require on every $u_{i}$ is that, for every two distributions $\delta,\delta^{\prime}\in\Delta(A)$:
$\begin{split}&\text{If }u_{i}(\delta^{\prime})>u_{i}(\delta),\\ &\text{then}\text{ for all }t\in(0,1),u_{i}(t\delta^{\prime}+(1-t)\delta)>u_{i}(\delta)\text{.}\end{split}$ (2)
Note that Condition (LABEL:eq:concavity) is implied by two stronger conditions:
  • Concavity — for all $t\in(0,1)$, $u_{i}(t\delta^{\prime}+(1-t)\delta)\geq tu_{i}(\delta^{\prime})+(1-t)u_{i}(\delta)$.
  • Strict quasiconcavity — for all $t\in(0,1)$, $u_{i}(t\delta^{\prime}+(1-t)\delta)>\min(u_{i}(\delta^{\prime}),u_{i}(\delta))$.
@@
Let $(N,A,(u_{i})_{i\in N},(w_{i})_{i\in N})$ be a weighted budget-aggregation game in which all $w_{i}$ are rational numbers and all $u_{i}$ satisfy Condition (LABEL:eq:concavity). Let $(\delta_{i^{c}})_{i^{c}\in N^{c}}$ be a Nash equilibrium of the corresponding cloned game. Define a strategy profile $(\delta_{i})_{i\in N}$ such that $\delta_{i}$ is the sum of $\delta_{i^{c}}$ for all $i^{c}$ which are clones of $i$. Then $(\delta_{i})_{i\in N}$ is a Nash equilibrium of the weighted game.
@@
Suppose that $(\delta_{i})_{i\in N}$ is not an NE, and let $i$ be a player with a profitable deviation $\delta^{\prime}_{i}$. Denote by $\delta,\delta^{\prime}$ the aggregated distribution before and after the deviation respectively (so $\delta^{\prime}=\delta-\delta_{i}+\delta^{\prime}_{i}$).
By definition of a profitable deviation, $u_{i}(\delta^{\prime})>u_{i}(\delta)$. By Condition (LABEL:eq:concavity), for every distribution $\delta^{\prime\prime}$ on the line between $\delta$ and $\delta^{\prime}$, $u_{i}(\delta^{\prime\prime})>u_{i}(\delta)$ holds too.
Take an arbitrary $\delta^{\prime\prime}\neq\delta$ on that line, that is sufficiently close to $\delta$ such that it can be attained by shifting an amount of $1/D$, that is, $\sum_{x\in A}|\delta^{\prime\prime}_{x}-\delta_{x}|\leq 2/D$. Let $\delta^{\prime\prime}_{i}:=\delta^{\prime\prime}-\delta+\delta_{i}$; note that $\sum_{x\in A}|\delta^{\prime\prime}_{i,x}-\delta_{i,x}|\leq 2/D$ too. Hence, it is possible to change $\delta_{i}$ to $\delta^{\prime\prime}_{i}$ by changing the strategy of one of the clones of $i$. As all clones have the same utility $u_{i}$, this implies that at least one of the clones has a profitable deviation.
Hence, the original profile is not an NE of the cloned game — a contradiction. ∎
@@
Given a class $C$ of utility functions satisfying Condition (LABEL:eq:concavity). If a NE can be computed in polynomial time for unweighted budget-aggregation games with utilities in class $C$, then a NE can be computed in pseudo-polynomial time for weighted budget-aggregation games with utilities in $C$.
@@
Given a weighted budget-aggregation game with weights $w_{i}=q_{i}/D$, construct the corresponding cloned game, which is an unweighted game. This game has $D$ agents; hence an NE can be computed in time polynomial in $m$ and $D$. By Lemma 8.2, this NE corresponds to a NE for the weighted game; the run-time is pseudo-polynomial in the parameters of that game. ∎
Let us check whether the utility functions considered in this paper satisfy Condition (LABEL:eq:concavity).
  • It is well-known that linear functions are concave, hence linear utilities satisfy (LABEL:eq:concavity).
  • It is also well-known that the minimum of concave functions is concave. As a Leontief utility function is defined as the minimum of (at most $|A|$) linear functions, it is concave and hence satisfies (LABEL:eq:concavity).
  • Binary symmetric separable utilities satisfy (LABEL:eq:concavity) if they are (strictly) concave, but not if they are strictly convex.
  • $\ell_{1}$ utilities are concave too. This holds more generally for any utility function derived from a norm-based metric $\|\cdot\|$, as for any $\delta,\delta^{\prime}$ and any $t\in(0,1)$,
    $u_{i}(t\delta^{\prime}+(1-t)\delta)$
    $=$ $-\|[t\delta^{\prime}+(1-t)\delta]-p_{i}\|$
    $=$ $-\|t(\delta^{\prime}-p_{i})+(1-t)(\delta-p_{i})\|$
    $\geq$ $-[\|t(\delta^{\prime}-p_{i})\|+\|(1-t)(\delta-p_{i})\|]$
        (by the triangle inequality)
    $=$ $-[t\|\delta^{\prime}-p_{i}\|+(1-t)\|\delta-p_{i}\|]$
        (by homogeneity of the norm)
    $=$ $tu_{i}(\delta^{\prime})+(1-t)u_{i}(\delta).$
    Hence $u_{i}$ is concave and satisfies (LABEL:eq:concavity).
Hence, for most classes of utility functions studied in this paper, Theorem 8.3 implies a pseudo-polytime algorithm for computing a NE in weighted budget-aggregation games. The case of binary symmetric separable strictly-convex utilities remains open. Moreover, it is open whether a NE in a weighted budget-aggregation game can be computed in time polynomial in $n$ and $m$, e.g., for Leontief utilities.
@@ 9 Conclusion
It would be interesting to see whether there exists a class of utility functions containing all of the previously considered preference models that still allow for an efficient computation of NEs. As a first step, one might want to extend binary symmetric separable utilities to allow for valuations other than 0,11 or consider preferences based on $\ell_{p}$ metrics other than $\ell_{1}$.

### v3

@@ 8 Weighted agents
In practice, one might want to assign different shares to the agents, e.g., when they represent groups of citizens of different sizes, or citizens in different tax brackets.
We define a weighted budget-aggregation game as a tuple $(N,A,\allowbreak(u_{i})_{i\in N},(w_{i})_{i\in N})$, where $N$, AA and $(u_{i})_{i\in N}$ are as in Definition 3.1, and $w_{i}$ is the weight of agent $i$, with $\sum_{i\in N}w_{i}=1$. The strategy set of each agent $i$ is given by
@@
For some classes of utility functions, the weighted setting can be reduced to the unweighted setting by agent cloning if all weights are rational numbers, resulting in a pseudo-polynomial time algorithm. We formalize the agent cloning process below.
@@
Given a weighted budget-aggregation game $(N,A,\allowbreak(u_{i})_{i\in N},(w_{i})_{i\in N})$ where all $w_{i}$ are rational numbers, let $D$ be their least common denominator, such that $w_{i}=q_{i}/D$ for all $i\in N$, and all $q_{i}$ are integers. The corresponding cloned game is an (unweighted) budget-aggregation game in which the set of agents $N^{c}$ contains $q_{i}$ clones of each agent $i$ (with the same utility function $u_{i}$), so that $|N^{c}|=\sum_{i\in N}q_{i}=D$.
@@
Let $(N,A,(u_{i})_{i\in N},(w_{i})_{i\in N})$ be a weighted budget-aggregation game in which all $w_{i}$ are rational numbers, and all $u_{i}$ are either linear, Leontief, binary symmetric separable and strictly concave/convex, or $\ell_{1}$. Let $(\delta_{i^{c}})_{i^{c}\in N^{c}}$ be a Nash equilibrium of the corresponding cloned game. Define a strategy profile $(\delta_{i})_{i\in N}$ such that $\delta_{i}$ is the sum of $\delta_{i^{c}}$ for all $i^{c}$ which are clones of $i$. Then $(\delta_{i})_{i\in N}$ is a Nash equilibrium of the weighted game.
@@
Suppose that $(\delta_{i})_{i\in N}$ is not an NE of the weighted game. For linear, Leontief, and $l_{1}$ preferences, there is a structural lemma that implies that, under $\delta$, some agent $i$ contributes a positive amount to some project $y$ in a set of “forbidden” projects that depends only on $u_{i}$ and $\delta$. Hence, under $\delta$, some clone $i_{c}$ of $i$ must contribute a positive amount to the same project $y$. As $i_{c}$ and $i$ have the same utility function, the set of “forbidden” projects under $\delta$ is the same for both of them; as $i_{c}$ contributes to a “forbidden” project, $\delta$ does not correspond to an NE of the cloned game.
Specifically:
  • For linear utilities, by Proposition 4.1, some agent $i$ contributes under $\delta$ to some non-highest-valued project $y\in A\setminus A^{\max}_{i}$;
  • For Leontief utilities, by Lemma 5.2, some agent $i$ contributes under $\delta$ to some non-critical project $y\in A\setminus T_{\delta,i}$;
  • For $\ell_{1}$ disutilities, by Lemma 7.1, some agent $i$ contributes under $\delta$ to some over-supplied project $y\in M^{>}_{i}(\delta)$.
It remains to prove the lemma for binary symmetric separable and strictly concave/convex preferences.
Theorem 6.2 remains true for the weighted game (Brandt et al., 2025a). Thus, some agent in the cloned game has to contribute to a non-critical alternative under $(\delta_{i})_{i\in N}$ in the cloned game and has a beneficial deviation, showing that the statement also holds for binary symmetric separable and strictly concave preferences.
Finally, if $(\delta_{i})_{i\in N}$ is no NE of the weighted game for binary symmetric separable and strictly convex preferences, this means that some agent $i$ contributes either to an alternative outside of $M_{\delta,i}$ or to two alternatives $x,y\in M_{\delta,i}$. For the first case, this implies that there exists a clone of agent $i^{c}$ that contributes to an alternative outside of $M_{\delta,i^{c}}$ in the cloned game. Corollary 6.5 proves that also for the second case, $(\delta_{i})_{i\in N}$ cannot refer to an NE of the cloned game.∎
@@
A Nash equilibrium can be computed in pseudo-polynomial time for weighted budget-aggregation games in which all utilities are either linear, Leontief, binary symmetric separable and strictly concave/convex, or $\ell_{1}$.
@@
Given a weighted budget-aggregation game with weights $w_{i}=q_{i}/D$, construct the corresponding cloned game, which is an unweighted game. This game has $D$ agents; hence, an NE can be computed in time polynomial in $m$ and $D$. By Lemma 8.2, this NE corresponds to an NE for the weighted game; the run-time is pseudo-polynomial in the parameters of that game. ∎
We believe Lemma 8.2 can be extended to other classes of utilities. It would be interesting to characterize the set of utility functions for which Lemma 8.2 remains true. Moreover, it is open whether an NE in a weighted budget-aggregation game can be computed in time polynomial in $n$ and $m$ (rather than pseudo-polynomial), e.g., for Leontief preferences.
@@ 9 Conclusion
It would be interesting to see whether there exists a class of utility functions containing all of the previously considered preference models that still allow for an efficient computation of NEs. As a first step, one might want to extend binary symmetric separable utilities to allow for valuations other than 0 or $1$, or consider preferences based on $\ell_{p}$ metrics other than $\ell_{1}$.
@@
Acknowledgments
We thank Yakov Babichenko, Felix Brandt, Warut Suksompong, and the anonymous reviewers of AAMAS 2026 for helpful comments. Patrick is supported by the Deutsche Forschungsgemeinschaft under grant BR 2312/14-1. Erel is supported by the Israel Science Foundation grant no. 1092/24.